OpenStudy (anonymous):

The excursion boat on the river takes 2½ hours to make the trip to a point 12 miles upstream and to return. If the rate at which the boat travels in still water is 5 times the rate of the river current, what is the rate of the current?The excursion boat on the river takes 2½ hours to make the trip to a point 12 miles upstream and to return. If the rate at which the boat travels in still water is 5 times the rate of the river current, what is the rate of the current?vvvv

OpenStudy (anonymous):

Can you help me robotobey?

OpenStudy (anonymous):

working on it.

OpenStudy (anonymous):

the answer is 2

OpenStudy (anonymous):

\[t=\frac{d}{r} \]Let c the current speed. Solve the following for c.\[\frac{12}{(5c-c)}+\frac{12}{(5c+c)}\text{=}\frac{5}{2} \]\[c=2 \text{ mph} \]

OpenStudy (anonymous):

Sorry it took so long. Had to deal with some local interruptions.